3. Find the relative extrema of f(x)= 2r3 +32- 12r-4, and use the second derivative test if applicable.
Answer #1

fz)= 2x3 +3- 12-4 1-X9 t9= ()t Critical Peiuts = 1-7-9 t279 72 17-2 22+2x--2 0 X42)-L(42) =0 (L42) (1-1)0 1=-2 1 ( -2,16) aud